OpenStudy (anonymous):

Determine whether the following polynomials can be factored by grouping. If so, factor the polynomials showing all work. If not, explain why this method will not work. A) 2c^2 + 4cd + 3c + 3d B) 4g^2 + 12gh + 5g + 15h

OpenStudy (zzr0ck3r):

2c^2+4cd+3c+3d= 2c(c+2d)+3(c+d) cant group 4g^2 +12gh +5g+15h= 4g(g+3h)+5(g+3h) = (4g+5)(g+3h) grouped

OpenStudy (anonymous):

thank you so much!

OpenStudy (zzr0ck3r):

note the first one is not because c+2d does not equal c+d

OpenStudy (anonymous):

okay

OpenStudy (zzr0ck3r):

or its not the same binomial

OpenStudy (anonymous):

okay